A compressor receives 10kW of power at 600 rpm from a motor. What is the torque acting on the shaft connecting the motor and the compressor?

System :Compressor receiving power from a motor through a shaft

To find : Torque


Known :. Power = 10kW

2. Speed = 600 rpm 1.

  


1
Expert's answer
2021-09-06T00:51:16-0400

Analysis


  1. w = 2"\pi nt"


Rate of work transferred = 10x103 KJ/s


speed

= 600 rpm

= 600 / 60 rps

= 10 rps


"Torque;"


"\therefore T=W/2\pi n"


"=10*1000/2\pi 10"


"=159.15N-m"